OpenStudy (anonymous):

5x − 2(x − 3) = 18 + x

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

5x-2x+6=18+x 6-18=x+2x-5x -12=-2x 12=2x 6=x